The hypotenuse of a right triangle measures 17 cm. One leg of the triangle measures 8 cm. Find the length of the other leg of the triangle.
Found 2 solutions by ewatrrr, MathTherapy:
Answer by ewatrrr(24785)


Hi

The hypotenuse of a right triangle measures 17 cm. 
One leg of the triangle measures 8 cm. 
Find the length of the other leg of the triangle.

Pythagorean Theorem:  
a^2 + b^2 = c^2
  a^2 + 8cm^2 = 17cm^2 
      a^2 = 225cm^2   Tossing out the negative solution for length
      c = 15 cm  

Important You feel comfortable using Your calculator
Wish You the Best in your Studies.

Answer by MathTherapy(10551)


The hypotenuse of a right triangle measures 17 cm. One leg of the triangle measures 8 cm. Find the length of the other leg of the triangle.
That's one of the few Pythagorean triples - the 8-15-17 "variety."